OpenStudy (anonymous):

Simplify. b−8 ______ b−4 Write your answer with a positive exponent only.

OpenStudy (anonymous):

b^-8 over b^-4

OpenStudy (anonymous):

how would I do this?

OpenStudy (anonymous):

@TheSmartOne can you help me again :( im like struggling over here

TheSmartOne (thesmartone):

Like this: \(\sf\Huge \frac{ b^{-8} }{b^{-4}}\)

OpenStudy (er.mohd.amir):

a^n/a^m= a^(n-m)

TheSmartOne (thesmartone):

Hint: \(\sf\LARGE\frac{a^m}{a^n}=a^{m-n}\) \(\sf\LARGE a^{-b}=\frac{1}{a^b}\)

OpenStudy (anonymous):



OpenStudy (anonymous):

I don't get that at all

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @TheSmartOne Hint: \(\sf\LARGE\frac{a^m}{a^n}=a^{m-n}\) \(\sf\LARGE a^{-b}=\frac{1}{a^b}\) \(\color{blue}{\text{End of Quote}}\) Let me change the variables. \(\sf\LARGE\frac{b^m}{b^n}=b^{m-n}\) \(\sf\LARGE b^{-c}=\frac{1}{b^c}\)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i got -8-(-4)= -4

TheSmartOne (thesmartone):

correct, now use the second formula :)